The operations manager of a large production plant would like to estimate the average amount of time workers take to assemble a new electronic component. After observing a number of workers assembling similar devices, she guesses that the standard deviation is 6 minutes. How large a sample of workers should she take if she wishes to estimate the mean assembly time to within 20 seconds? Assume that the confidence level is to be 99%.
1
Expert's answer
2020-06-08T19:25:38-0400

"P\{|x^{-}|=\frac{\sigma}{\sqrt {n}}x\}=2\Phi(x)\\
x^{-}=20=\frac{20}{60}=\frac{1}{3}\\
\sigma=6\\
2\Phi(x)=0.99\\
\Phi(x)=0.495\\
x=2.58\\
\frac{1}{3}=\frac{6}{\sqrt {n}}\cdot2.58\\
\sqrt {n}=46.44\\
n=2156,6736\\
n=2157"